OpenStudy (anonymous):

DO NOT JUST GIVE THE ANSWER. HELP WORK IT OUT. What is the ratio of the surface area of a sphere with radius 2 ft to the surface area of a sphere with radius 5 ft?

OpenStudy (anonymous):

what is the formula for surface area of a sphere?

OpenStudy (anonymous):

@Cal3blikespie

OpenStudy (anonymous):

A = 4 * π * r^2

OpenStudy (anonymous):

good so, \[\frac{ 4*\pi * 2^{2} }{ 4*\pi * 5^{2} }\]

OpenStudy (anonymous):

The 4 and pi cancel out

OpenStudy (anonymous):

We have 2^2/5^2 4/25

OpenStudy (anonymous):

0.16?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

4/25 is 0.16

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and so is 2^2/5^2

OpenStudy (anonymous):

so now what?

OpenStudy (anonymous):

is the ratio 2 to 5?

OpenStudy (anonymous):

4 to 25

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

We have 2^2/5^2 4/25

OpenStudy (anonymous):

I solved it see from the begining

OpenStudy (anonymous):

4∗π∗224∗π∗52

OpenStudy (anonymous):

ugh. you said the four and pi cancel out

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

OMG

OpenStudy (anonymous):

i understand how you got 5^2 and 2^2

OpenStudy (anonymous):

but where did the 4/25 come from?

OpenStudy (anonymous):

So?

OpenStudy (anonymous):

2^2 is 4 5^2 is 25 So, 4/25

OpenStudy (anonymous):

= 4:25

OpenStudy (anonymous):

oooooooh. ok i get it

OpenStudy (anonymous):

so when the ratio would be 4 to 25

OpenStudy (anonymous):

then

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Welcome